OpenStudy (anonymous):

2sin3x=sqrt3 interval (0,12pi)

OpenStudy (campbell_st):

well start with \[\sin(3x) = \frac{\sqrt{3}}{2}\] which is an exact value and is in the 1st and 2nd quadrant.. find the angle... then divide by 3